OpenStudy (anonymous):

set solution of -[8z-(15z+10)]=10+(6z+4)please

OpenStudy (anonymous):

Bump your question:P

OpenStudy (anonymous):

is not letting me..yet);

OpenStudy (anonymous):

how many min ar left? once it lets you bump it kk:]

OpenStudy (anonymous):

k..got it(: it says 3 mins

OpenStudy (anonymous):

kk let me know when you did it kk:]

OpenStudy (anonymous):

k..i'm about to bump it!(:

OpenStudy (anonymous):

bumped!

OpenStudy (anonymous):

kk cool

OpenStudy (anonymous):

@BrennaM @evergirl @Falco276 can you please help with this problem please:]

OpenStudy (anonymous):

hey @ryan123345

OpenStudy (ryan123345):

hi :P

OpenStudy (anonymous):

can u help with problem:P

OpenStudy (anonymous):

@ryan123345

OpenStudy (anonymous):

pls or find someone tht can help?

OpenStudy (anonymous):

tag @hero he's a math wiz :P

OpenStudy (anonymous):

lol thnx girl:P @Hero

OpenStudy (anonymous):

arnt you suppost to use pemdas?

OpenStudy (ryan123345):

u habe the possible answers? i just solved it.

OpenStudy (anonymous):

please refresh my memory inverse opper.?

OpenStudy (anonymous):

this is what i ended up with 12t-30=60-6t

OpenStudy (anonymous):

and then i got lost

OpenStudy (anonymous):

- 30 and 60

OpenStudy (anonymous):

+30 imean

OpenStudy (anonymous):

sorry wrong problem..

OpenStudy (anonymous):

oh well for tht one u would get 15 i belive:P

OpenStudy (anonymous):

this one is the correct one -11z+2=3z+12

OpenStudy (anonymous):

thats what i ended up with

OpenStudy (anonymous):

you have to -2 frm 12

OpenStudy (falco276):

cant do that, sorry :(

OpenStudy (anonymous):

which gives me -11z=3z+10

OpenStudy (anonymous):

Don't feel bad, I couldn't have even gotten as far as u did xD

OpenStudy (anonymous):

omg ibarra :P -11z and 3z

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

ugh..gotta get ready for class..ill be back later..thank you All soo much(:

OpenStudy (anonymous):

-14z=10 divide:P